OpenStudy (anonymous):

Solve the following equation in the complex number system. Show work.. x^5+20x^3-x^2-20=0

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

that's what I got , just checking.. thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

It should have 5 solutions

OpenStudy (anonymous):

\[x^5+20x^3-x^2-20=x^3(x^2+20)-(x^2+20)\]\[=(x^3-1)(x^2+20)=(x-1)(x^2+x+1)(x^2+20)=0\]So you need the complex solutions to:\[x^2+x+1=0\]and:\[x^2+20=0\]which can be obtained by quadratic formula or whatever.

OpenStudy (anonymous):

Some factoring by grouping right there. Nice

OpenStudy (anonymous):

Thanks mr. joe math...

OpenStudy (anonymous):

nice factoring!

OpenStudy (anonymous):

it was a setup, im promise! lol

OpenStudy (anonymous):

i knew that!